If f(x) = 2x^2 - 1 and h(x) = 18x^4 - 24x^2 + 7, determine a formula for g(x) such that f ∘ g = h
Answer by stanbon(75887)

If f(x) = 2x^2 - 1 and h(x) = 18x^4 - 24x^2 + 7, determine a formula for g(x) such that f ∘ g = h
-------
f[g(x)] = h(x)
----
2(g(x))^2 -1 = 18x^4 - 24x^2 + 7
----------------------
g(x)^2 = [18x^4-24x^2+7 + 1]/2 = 9x^4 - 12x^2 + 4
------------------
Take the square root to get:
g(x) = (3x^2 - 2)
-------------------------
Cheers,
Stan H.